# Bounds of Validity for Bifurcations of Equilibria in a Class of Networked Dynamical Systems

### v2

@@ IV-B Firing Rate models
Define projection matrices $P:=P_{\operatorname{range}(J^{\star}_{F})}=W_{F}W_{F}^{\top}$ and $Q:=P_{\ker(J^{\star}_{F})^{\perp}}=\bar{V}_{F}\bar{V}_{F}^{\top}$ as described in (1), (2) and (3). We make the following assumption:
@@
    The pair $(x^{\star},p^{\star})$ is an equilibrium of (13), i.e.,
    $\Psi(x^{\star},p^{\star})=-Cx^{\star}+S(p^{\star}Ax^{\star}+b)=0.$ (22)
@@
    The Jacobian $J^{\star}_{F}$ has a zero eigenvalue with algebraic and geometric multiplicity $q$, and $\mathrm{Re}\,(\lambda_{j})\neq 0$ for all other eigenvalues.
@@
$\left\lVert{\xi_{\parallel}(\alpha,p)}\right\rVert=\left\lVert{P\begin{bmatrix}\Delta^{\parallel}_{x}(\alpha,p)AV&\Delta^{\parallel}_{p}(\alpha,p)\end{bmatrix}}\right\rVert$ where
@@
$\hskip 91.04881pt-D_{z}S(p^{\star}Ax^{\star}+b)Ax^{\star}$ (23b)
@@
Following the definition (7a), we must compute $\xi_{\parallel\alpha}(\alpha,p)$ and $\xi_{\parallel p}(\alpha,p)$. Since $J^{\star}_{F}V=0$ we can write
$CV=p^{\star}D_{z}S(p^{\star}Ax^{\star}+b)AV$
which allows us to rewrite $\xi_{\parallel\alpha}(\alpha,p)$ as
@@
$\hskip 25.60747pt-p^{\star}D_{z}S(p^{\star}Ax^{\star}+b)\bigr]AV=:\Delta^{\parallel}_{x}(\alpha,p)AV$
Next, we rewrite $\xi_{\parallel p}(\alpha,p)$ as
$\xi_{\parallel p}(\alpha,p)$ $=D_{p}\Psi(V\alpha+\bar{V}\beta_{0},p)-D_{p}\Psi(x^{\star},p^{\star})$
@@
$\hskip 45.5244pt-D_{z}S(p^{\star}Ax^{\star}+b)Ax^{\star}\bigr]=:\Delta^{\parallel}_{p}(\alpha,p)$
Finally, combining our expressions for $\xi_{\parallel\alpha}$ and $\xi_{\parallel p}$ yields
$\xi_{\parallel}(\alpha,p)=W^{\top}\begin{bmatrix}\Delta^{\parallel}_{x}(\alpha,p)AV&\Delta^{\parallel}_{p}(\alpha,p)\end{bmatrix}$.
The desired expression for $\left\lVert{\xi_{\parallel}(\alpha,p)}\right\rVert$ follows directly by applying II.1 with $Y=A\begin{bmatrix}\Delta^{\parallel}_{x}(\alpha,p)V&\Delta^{\parallel}_{p}(\alpha)\end{bmatrix}$. ∎
@@
$\Delta_{\perp}(\alpha,\beta,p):=pD_{z}S(pAx+b)-p^{\star}D_{z}S(p^{\star}Ax^{\star}+b)$ (24)
@@
$\xi_{\perp}(\alpha,\beta,p)$ $=W^{\top}\big(D_{x}\Psi(\Gamma(\alpha,\beta),p)-D_{x}(x^{\star},p^{\star})\big)\bar{V}$
$=W^{\top}\Bigl(-C+pD_{z}S\bigl(pA\Gamma(\alpha,\beta)+b\bigr)$
$\hskip 42.67912pt+C-p^{\star}D_{z}S(p^{\star}Ax^{\star}+b)\Bigr)A\bar{V}$
$=W^{\top}\Delta_{\perp}(\alpha,\beta,p)A\bar{V}.$
Applying II.1 and then II.2 to the above results in the desired expression for $\xi_{\perp}(\alpha,\beta,p)$. ∎
Finally, we state an expression of the bounds for (13).
Theorem IV.6 (Bounds of bifurcation in Firing Rate models)
Consider (13) under 2. For all $r_{\parallel},r_{\perp}>0$ satisfying
@@
where
$L_{\parallel}(r_{\parallel})$ $=\sup\Bigl\{\left\lVert{P\begin{bmatrix}\Delta^{\parallel}_{x}(\alpha,p)AV&\Delta^{\parallel}_{p}(\alpha,p)\end{bmatrix}}\right\rVert$
$\hskip 51.21495pt:(\alpha,p)\in\mathcal{B}\bigl((\alpha_{0},p^{\star}),r_{\parallel}\bigr)\Bigr\}$
$L_{\perp}(r_{\parallel},r_{\perp})$ $=\sup\Bigl\{\left\lVert{P\Delta_{\perp}(\alpha,\beta,p)AQ}\right\rVert$
$:(\alpha,p)\in\mathcal{B}\bigl((\alpha_{0},p^{\star}),r_{\parallel}\bigr),\;\beta\in\mathcal{B}(\beta_{0},r_{\perp})\Bigr\},$
with $\Delta_{x}^{\parallel}$, $\Delta_{p}^{\parallel}$, $\Delta_{\perp}$ are defined in (23), (24), there exists a smooth implicit map $\varphi:\mathcal{B}\bigl((\alpha_{0},p^{\star}),r_{\parallel}\bigr)\to\mathcal{B}(\beta_{0},r_{\perp})$ that satisfies (5).
@@
Substituting the values of (6a), (6b), (8a), (8b) from III.1, III.2, IV.4, IV.5 into (10) from Theorem III.3 yields the theorem statement. ∎
@@ V Application to Consensus Bifurcations
In recent years, models of the form (12), (13) have received attention in the literature on collective decision-making and opinion dynamics. In this context, components of the state vector $x$ represent scalar opinions of individuals in a social network, which are updated in time according to a nonlinear saturating social rule. Of particular interest is the emergence of collective decision equilibria as bifurcations from the neutral $x=0$ state. Such indecision-breaking bifurcations have been established in opinion dynamics models with both Hopfield-like structure [franci2015realization, gray2018multiagent, fontan2017multiequilibria] and firing rate-like structure [bizyaeva2021patterns, bizyaeva2023nonlinear, leonard2024fast]. In this section, we apply the bounds established in section IV to assess the validity region of the Lyapunov–Schmidt reduction analysis of indecision-breaking supercritical pitchfork bifurcations in these nonlinear opinion dynamic models over regular balanced consensus graphs.
To specialize (12), (13) to nonlinear opinion dynamics (NOD) over regular graphs, we set $C=dI_{n}$ with $d>0$, $p=u$, let $A$ be the adjacency matrix of a $k$-regular, undirected and connected simple graph with unit edge weights, $S=\tanh$, and consider the unbiased case $b\equiv\mathbf{0}_{n}$. By the definition of a $k$-regular graph, the adjacency matrix $A$ satisfies $A\mathbf{1}_{n}=k\mathbf{1}_{n}$. For both models, there is an indecision-breaking bifurcation point $(x^{\star},u^{\star})=\left(\mathbf{0}_{n},\frac{d}{k}\right)$ at which consensus solutions emerge, see Fig. 1 for schematic. Note from (4) that $x^{\star}=\mathbf{0}_{n}\implies(\alpha_{0},\beta_{0})=(0,\mathbf{0}_{n-1})$.
@@ V-A Hopfield-structured NOD
We first characterize the orthonormal basis matrix $V$ from (2). By the Perron–Frobenius theorem [horn2012matrix, Theorem 8.4.4], the adjacency matrix $A$ of a $k$-regular, irreducible graph admits a simple largest eigenvalue $k$, with associated Perron eigenvector $\mathbf{1}_{n}$. Consequently, $\ker(J)$ is one-dimensional and spanned by the normalized vector $V=\tfrac{1}{\sqrt{n}}\mathbf{1}_{n}$. Also note that since $J^{\star}$ is symmetric, we have $W=\bar{V}$ and $V=\bar{W}$, and so $P=Q$ and $VV^{\top}+WW^{\top}=I_{n}$.
@@
$M_{\parallel}=\tfrac{1}{u^{\star}}\left\lVert{P(Cx^{\star}-b)}\right\rVert=\tfrac{k}{d}\left\lVert{(I-VV^{\top})(dI_{n}\mathbf{0}_{n})}\right\rVert=0.$
To evaluate $L_{\parallel}(r_{\parallel})$, we first evaluate the terms $\Delta_{\parallel}(\alpha,u)$ and $\bar{S}_{\parallel}(\alpha)$ defined in (16) for $\xi_{\parallel}(\alpha,u)$ to obtain
@@
Substituting these expressions into IV.1 yields
@@
and hence, the smallest nonzero singular value of $J^{\star}$ is $\lambda_{2}(J^{\star})=\tfrac{d}{k}(\lambda_{2}(A)-k)$ Therefore,
@@
Now, to evaluate $L_{\perp}(r_{\parallel},r_{\perp})$, we first compute the expression for $\Delta_{\perp}(\alpha,\beta,u)$ defined in (16) for IV.2 as
@@
Having computed all requisite quantities, we now state the bounds for the Hopfield regular balanced consensus network:
@@
The bounds of validity of the Lyapunov–Schmidt reduction from II.3 around the pitchfork bifurcation at $(\mathbf{x}^{\star},u^{\star})=(\mathbf{0},\tfrac{d}{k})$ for the $n$-node $k$-regular balanced consensus opinion dynamic network (27) is
@@
Substituting the values of (6b) and $M_{\parallel}=0$, $L_{\parallel}(r_{\parallel})=0$, $M_{\perp}=k/(d(k-\lambda_{2}(A))$, $L_{\perp}(r_{\parallel},r_{\perp})=r_{\perp}|\lambda^{\prime}(A)|$ into the inequality (10) from Theorem III.3 yields the desired bounds. ∎
@@ V-B Firing-Rate-structured NOD
where $s_{uk\alpha}=\operatorname{sech}^{2}(uk\alpha/\sqrt{n})$. Substituting into IV.4
@@
since both terms both terms lie in the span of $V$, which is annihilated by $Q$. From (8a) we conclude that $L_{\parallel}(r_{\parallel})=0$. Again, inequality (10) reduces to $M_{\perp}L_{\perp}(r_{\parallel},r_{\perp})<1$. It remains to compute $M_{\perp}$ and $L_{\perp}(r_{\parallel},r_{\perp})$.
Since $J^{\star}$ remains unchanged from the Hopfield model, we identically have $M_{\perp}=\frac{k}{d\big(k-\lambda_{2}(A)\big)}$. To justify that $L_{\perp}(r_{\parallel},r_{\perp})$ also remains unchanged, we compute the expression for defined in (24) for IV.5 to get $\Delta_{\perp}(\alpha,\beta,u)=u\operatorname{diag}\bigl(\operatorname{sech}^{2}\bigl(uA\Gamma(\alpha,\beta)\bigr)\bigr)-\tfrac{d}{k}\operatorname{diag}\bigl(\operatorname{sech}^{2}(\tfrac{d}{k}A\mathbf{0}_{n})\bigr)=u\operatorname{diag}\bigl(\operatorname{sech}^{2}\bigl(uA\Gamma(\alpha,\beta)\bigr)\bigr)-\tfrac{d}{k}I_{n}.$ By the same reasoning as in the Hopfield case, we can say that that $\left\lVert{\xi_{\parallel}(\alpha,u)}\right\rVert$ attains its supremum at $(\alpha,\beta,u)=(0,\mathbf{0}_{n-1},u^{\star}\pm r_{\parallel})$. Consequently, identical to the Hopfield case, $L_{\perp}(r_{\parallel},r_{\perp})=r_{\parallel}\left\lvert{\lambda^{\prime}(A)}\right\rvert$. Therefore, the bounds of validity for the reduced order model obtained via Lyapunov–Schmidt reduction at the indecision-breaking consensus bifurcation in the Firing Rate-form NOD model (29) coincide with those of the Hopfield-form NOD model (27), as stated in Theorem V.1. We illustrate how these bounds depend on the network size $n$ and node degree $k$ in Figure 2.
@@ VI Discussion
In this article we established explicit, computable bounds for the validity of Lyapunov–Schmidt reduction-based bifurcation analysis in representative classes of networked nonlinear systems, and demonstrated their utility by applying the bounds to an illustrative class of nonlinear consensus problems on regular graphs, for which the evaluation proves to be analytically computable. Notably, the derived expressions depend directly on interpretable features, such as the spectral properties of the underlying graph, providing insight into how network structure influences these bounds. In future work, we will aim to identify additional classes of systems with bifurcations analyzed with the Lyapunov–Schmidt reduction for which these bounds can be computed analytically.

### v3

@@ IV-B Firing Rate models
Define projection matrices $P:=P_{\operatorname{range}(J^{\star}_{F})}=W_{F}W_{F}^{\top}$ and $Q:=P_{\ker(J^{\star}_{F})^{\perp}}=\overline{V}_{F}\overline{V}_{F}^{\top}$ as described in (1) and (2). We make the following assumption:
@@
    The pair $(x^{\star},p^{\star})$ is an equilibrium of (11).
@@
    The Jacobian $J^{\star}_{F}$ has a zero eigenvalue with multiplicity $q$, and $\mathrm{Re}\,(\lambda_{j})\neq 0$ for all other eigenvalues.
@@
$\left\lVert{\xi_{\parallel}(\alpha,p)}\right\rVert=\left\lVert{P\begin{bmatrix}\Delta^{\parallel}_{x}(\alpha,p)AV&\Delta^{\parallel}_{p}(\alpha,p)\end{bmatrix}}\right\rVert$ where $\Delta^{\parallel}_{x}(\alpha,p)$ and $\Delta^{\parallel}_{p}(\alpha,p)$ are defined as
@@
$\hskip 62.59605pt-D_{z}S(p^{\star}Ax^{\star}+b)Ax^{\star}$ (18b)
@@
Following the definition (6a), we must compute $\xi_{\parallel\alpha}(\alpha,p)$ and $\xi_{\parallel p}(\alpha,p)$. Since $J^{\star}_{F}V=0$ we can write $CV=p^{\star}D_{z}S(p^{\star}Ax^{\star}+b)AV,$ simplifying $\xi_{\parallel\alpha}(\alpha,p)$ as
@@
$\hskip 5.69054pt-p^{\star}D_{z}S(p^{\star}Ax^{\star}+b)\bigr]AV=:\Delta^{\parallel}_{x}(\alpha,p)AV$
Next, we express
$\xi_{\parallel p}(\alpha,p)$ $=D_{p}\Psi(\Gamma(\alpha,\beta_{0}),p)-D_{p}\Psi(x^{\star},p^{\star})$
@@
$\hskip 28.45274pt-D_{z}S(p^{\star}Ax^{\star}+b)Ax^{\star}\bigr]=:\Delta^{\parallel}_{p}(\alpha,p)$
Finally, combining our expressions for $\xi_{\parallel\alpha}$ and $\xi_{\parallel p}$ yields $\xi_{\parallel}(\alpha,p)=W^{\top}\begin{bmatrix}\Delta^{\parallel}_{x}(\alpha,p)AV&\Delta^{\parallel}_{p}(\alpha,p)\end{bmatrix}$. Applying Lemma II.1 to this yields the desired expression. ∎
@@
$\Delta_{\perp}(\alpha,\beta,p):=pD_{z}S(pAx+b)-p^{\star}D_{z}S(p^{\star}Ax^{\star}+b)\vskip 5.69054pt$ (19)
@@
$\xi_{\perp}(\alpha,\beta,p)$ $=W^{\top}\big(D_{x}\Psi(\Gamma(\alpha,\beta),p)-D_{x}\Phi(x^{\star},p^{\star})\big)\overline{V}$
$=W^{\top}\bigl(pD_{z}S\bigl(pA\Gamma(\alpha,\beta)+b\bigr)$
$\hskip 36.98857pt-p^{\star}D_{z}S(p^{\star}Ax^{\star}+b)\bigr)A\overline{V}$
$=W^{\top}\Delta_{\perp}(\alpha,\beta,p)A\overline{V}.$
Applying Lemma II.1 to the above results in the desired expression for $\xi_{\perp}(\alpha,\beta,p)$, since $\left\lVert{W^{\top}\Delta_{\perp}(\alpha,\beta,p)A\overline{V}}\right\rVert=\left\lVert{WW^{\top}\Delta_{\perp}(\alpha,\beta,p)A\overline{V}\,\overline{V}^{\top}}\right\rVert=\left\lVert{P\Delta_{\perp}(\alpha,\beta,p)AQ}\right\rVert.$
With these lemmas established, we now state the bounds from Theorem III.3 for bifurcations in the Firing Rate model.
Theorem IV.6
Consider the Firing Rate dynamics (11) under 2. For all $r_{\parallel},r_{\perp}>0$ satisfying
@@
$L_{\parallel}(r_{\parallel})$ $=\underset{(\alpha,p)\in\mathcal{B}_{\parallel}(r_{\parallel})}{\sup}\left\lVert{P\begin{bmatrix}\Delta^{\parallel}_{x}(\alpha,p)AV&\Delta^{\parallel}_{p}(\alpha,p)\end{bmatrix}}\right\rVert$
$L_{\perp}(r_{\parallel},r_{\perp})$ $=\underset{\begin{subarray}{c}(\alpha,p)\in\mathcal{B}_{\parallel}(r_{\parallel})\\ \beta\in\mathcal{B}_{\perp}(r_{\perp})\end{subarray}}{\sup}\left\lVert{P\Delta_{\perp}(\alpha,\beta,p)AQ}\right\rVert$
with $\Delta_{x}^{\parallel}$, $\Delta_{p}^{\parallel}$, $\Delta_{\perp}$ defined in (18), (19), there exists a smooth implicit map $\varphi:\mathcal{B}_{\parallel}(r_{\parallel})\to\mathcal{B}_{\perp}(r_{\perp})$ that satisfies (4).
@@
Substituting (5b), (7a), (7b) from Lemmas III.2, IV.4, IV.5 into (9) from Theorem III.3 yields the theorem statement. ∎
@@ V Application to Consensus Bifurcations
In recent years, models of the form (10), (11) have received attention in the literature on collective decision-making and opinion dynamics. In this context, components of the state vector $x$ represent scalar opinions of individuals in a social network, which are updated in time according to a nonlinear saturating social rule. Of particular interest is the emergence of collective decision equilibria as bifurcations from the neutral $x=0$ state. Such indecision-breaking bifurcations have been established in opinion dynamics models with both Hopfield-like structure [7, 12, 6] and firing rate-like structure [5, 4, 17]. In this section, we apply the bounds established in Section IV to assess the validity region of the Lyapunov–Schmidt reduction-based analysis of indecision-breaking supercritical pitchfork bifurcations in these nonlinear opinion dynamic models over regular balanced consensus graphs.
To specialize (10), (11) to nonlinear opinion dynamics (NOD) over regular graphs, we set $C=dI_{n}$ with $d>0$, $p=u$, let $A$ be the adjacency matrix of a $k$-regular, undirected and connected simple graph with unit edge weights, $S=\tanh$, and consider the unbiased case $b\equiv\mathbf{0}_{n}$. By the definition of a $k$-regular graph, the adjacency matrix $A$ satisfies $A\mathbf{1}_{n}=k\mathbf{1}_{n}$. For both models, there is an indecision-breaking bifurcation point $(x^{\star},u^{\star})=\left(\mathbf{0}_{n},\frac{d}{k}\right)$ at which consensus solutions emerge, see Figure 1 for schematic. Note from (3) that $x^{\star}=\mathbf{0}_{n}\implies(\alpha_{0},\beta_{0})=(0,\mathbf{0}_{n-1})$.
@@ V-A Hopfield-structured NOD
We first characterize the orthonormal basis matrix $V$ of $\ker(J^{\star})$. By the Perron–Frobenius theorem [15, Theorem 8.4.4], the adjacency matrix $A$ of a $k$-regular, irreducible graph admits a simple largest eigenvalue $k$, with associated eigenvector $\mathbf{1}_{n}$. Consequently, $\ker(J^{\star})$ is one-dimensional and spanned by the normalized vector $V=\tfrac{1}{\sqrt{n}}\mathbf{1}_{n}$. Also note that since $J^{\star}$ is symmetric, we have $W=\overline{V}$ and $V=\overline{W}$, and so $P=Q$ and $VV^{\top}+WW^{\top}=I_{n}$.
@@
$M_{\parallel}=\tfrac{1}{\left\lvert{u^{\star}}\right\rvert}\left\lVert{P(Cx^{\star}-b)}\right\rVert=\tfrac{k}{d}\left\lVert{(I-VV^{\top})(dI_{n}\mathbf{0}_{n})}\right\rVert=0.$
To evaluate $L_{\parallel}(r_{\parallel})$, we first evaluate the terms $\Delta_{\parallel}(\alpha,u)$ and $\bar{S}_{\parallel}(\alpha)$ defined in (13a) for $\xi_{\parallel}(\alpha,u)$ to obtain
@@
Substituting these expressions into Lemma IV.1 yields
@@
and hence, the smallest non-zero singular value of $J^{\star}$ is $\lambda_{2}(J^{\star})=\tfrac{d}{k}(\lambda_{2}(A)-k)$ Therefore,
@@
Now, to evaluate $L_{\perp}(r_{\parallel},r_{\perp})$, we first compute the expression for $\Delta_{\perp}(\alpha,\beta,u)$ defined in (13a) for Lemma IV.2 as
@@
Having computed all requisite quantities, we now state the bounds for the Hopfield regular balanced consensus network.
@@
The bounds of validity of the Lyapunov–Schmidt reduction from Theorem II.2 around the pitchfork bifurcation at $(\mathbf{x}^{\star},u^{\star})=(\mathbf{0},\tfrac{d}{k})$ for the $n$-node $k$-regular balanced consensus opinion dynamic network (21) is
@@
Substituting the values of (5b) and $M_{\parallel}=0$, $L_{\parallel}(r_{\parallel})=0$, $M_{\perp}=k/(d(k-\lambda_{2}(A))$, $L_{\perp}(r_{\parallel},r_{\perp})=r_{\parallel}|\lambda^{\prime}(A)|$ into the inequality (9) from Theorem III.3 yields the desired bounds. ∎
@@ V-B Firing-Rate-structured NOD
where $s_{uk\alpha}=\operatorname{sech}^{2}(uk\alpha/\sqrt{n})$. Substituting into Lemma IV.4
@@
since both terms lie in the span of $V$, which is annihilated by $Q$. From (7a) we conclude that $L_{\parallel}(r_{\parallel})=0$. Again, inequality (9) reduces to $M_{\perp}L_{\perp}(r_{\parallel},r_{\perp})<1$. It remains to compute $M_{\perp}$ and $L_{\perp}(r_{\parallel},r_{\perp})$.
Since $J^{\star}$ remains unchanged from the Hopfield model, we identically have $M_{\perp}=\frac{k}{d\big(k-\lambda_{2}(A)\big)}$. To justify that $L_{\perp}(r_{\parallel},r_{\perp})$ also remains unchanged, we compute the expression for $\Delta_{\perp}$ defined in (19) for Lemma IV.5 to get
$\Delta_{\perp}(\alpha,\beta,u)$ $=u\operatorname{diag}\bigl(\operatorname{sech}^{2}\bigl(uA\Gamma(\alpha,\beta)\bigr)\bigr)$
$\qquad-\tfrac{d}{k}\operatorname{diag}\bigl(\operatorname{sech}^{2}(\tfrac{d}{k}A\mathbf{0}_{n})\bigr)$
$=u\operatorname{diag}\bigl(\operatorname{sech}^{2}\bigl(uA\Gamma(\alpha,\beta)\bigr)\bigr)-\tfrac{d}{k}I_{n}.$
With the same reasoning as the Hopfield case, we can say that $\left\lVert{\xi_{\parallel}(\alpha,u)}\right\rVert$ attains its supremum at $(\alpha,\beta,u)=(0,\mathbf{0}_{n-1},u^{\star}\pm r_{\parallel})$. Consequently, identical to the Hopfield case, $L_{\perp}(r_{\parallel},r_{\perp})=r_{\parallel}\left\lvert{\lambda^{\prime}(A)}\right\rvert$. Therefore, the bounds of validity for the reduced order bifurcation equations obtained via Lyapunov–Schmidt reduction at the indecision-breaking bifurcation in the Firing Rate-form NOD model (23) coincide with those of the Hopfield-form NOD model (21), as stated in Theorem V.1. We illustrate how these bounds depend on the network size $n$ and node degree $k$ in Figure 2.
@@ VI Discussion
In this article we established explicit, computable bounds for the validity of Lyapunov–Schmidt reduction-based bifurcation analysis in representative classes of networked nonlinear systems, and demonstrated their utility by applying the bounds to an illustrative class of nonlinear consensus problems on regular graphs, for which the evaluation proves to be analytically computable. Notably, the derived expressions depend directly on interpretable features, such as the spectral properties of the underlying graph, providing insight into how network structure influences these bounds. In future work, we will aim to identify additional classes of systems with bifurcations analysed with the Lyapunov–Schmidt reduction for which these bounds can be computed analytically, and to analyse how these bounds behave across other graph families, such as Erdős-Rényi and small-world networks.
VII Acknowledgement
Ravi N Banavar’s work was partially supported by a MATRICS grant from the Anusandhan Research Foundation (ANRF).